OpenStudy (anonymous):

My head hurts! Any math loving folks able to shine a light on this problem? Simplify (sqrt48)-6 10 Pretend this is a fraction (not sure how to enter all that in). So I'm not sure if I should - the 6 first or simplify the sqrt. I decided the square should be simplified first but its an imperfect square. I am temporarily giving up and admitting I need help. :/

OpenStudy (mokeira):

\[\sqrt{46}-6\] Is that the question?

OpenStudy (anonymous):

46 should be 48 and then what you have is over 10. :)

OpenStudy (mokeira):

or is it \[\frac{ \sqrt{46}-6 }{ 10 }\]

OpenStudy (anonymous):

Thats right except for the 46 needs to be 48. Or is my screen just goofing up?

OpenStudy (mokeira):

lol..its my mistake..

OpenStudy (mokeira):

\[\frac{ \sqrt{48}-6 }{ 10 }\] square everything \[(\frac{ \sqrt{48}-6 }{10 })^{2}\] \[\frac{ 48-36 }{ 100 }\] can you go on from there?

OpenStudy (anonymous):

Silly question... is it standard in math to do what you did? Like a no brainer, next step sort of thing?

OpenStudy (anonymous):

I know I do that when I complete the square of a quadratic equation, but I never thought I had permission to do that for another problem.

OpenStudy (mokeira):

haha lol...math is all about being creative. Using the basics you have

OpenStudy (anonymous):

Some part of my human nature is doubting it is really that simple. Thanks for your feedback :)

OpenStudy (mokeira):

you welcome!

OpenStudy (anonymous):

I went on math site to check this answer and it comes up with something completely different. Any idea why?

OpenStudy (mokeira):

what does it come with

OpenStudy (mokeira):

I mean what is the answer

OpenStudy (anonymous):

ok let me write it down

OpenStudy (anonymous):

Woah my life just got changed. I just saw the equation button! \[\frac{ 1 }{ 5} (2\sqrt{3} -3)\]

OpenStudy (anonymous):

There it is :D

OpenStudy (anonymous):

But sometimes I don't trust mathway, so don't please dont be offended by my question ;)

OpenStudy (mokeira):

this is how they did it \[\sqrt{48}\] is the same as \[\sqrt{16\times3}\] so \[\frac{ \sqrt{16\times3}-6 }{ 10 }\] \[\frac{ 4\sqrt{3}-6 }{ 10 }\] Then simplified by dividing by 2 \[\frac{ 2\sqrt{3} -3}{ 5 }\] which is same as \[\frac{ 1 }{ 5 }(2\sqrt{3}-3)\]

OpenStudy (mokeira):

lol..thats how

OpenStudy (anonymous):

good gravy. I feel overwhelmed. How do you know which answer is right?

OpenStudy (anonymous):

AWESOME explanation by the way :D

OpenStudy (mokeira):

i was wrong :(

OpenStudy (mokeira):

mathway was right

OpenStudy (anonymous):

Is there anything I can do to give you a medal? I'm blown away from the time you took to help me and your humbleness too

OpenStudy (anonymous):

(but I'm not familiar with this site so not sure how to do that)

OpenStudy (mokeira):

hahaha lol..no medal needed..

OpenStudy (mokeira):

im here to learn too

OpenStudy (anonymous):

I'm always so curious about the people who answer these. So far you are my favorite. Thanks a million, Mokeira!

OpenStudy (mokeira):

lol @LaylaJ ...just practise they are not hard

OpenStudy (mokeira):

Glad to help